# Deep Learning Approach for Changepoint Detection: Penalty Parameter Optimization

### v2

@@ Abstract
Changepoint detection, a technique for identifying significant shifts within data sequences, is crucial in various fields such as finance, genomics, medicine, etc. Dynamic programming changepoint detection algorithms are employed to identify the locations of changepoints within a sequence, which rely on a penalty parameter to regulate the number of changepoints. To estimate this penalty parameter, previous work uses simple models such as linear models or decision trees. This study introduces a novel deep learning method for predicting penalty parameters, leading to demonstrably improved changepoint detection accuracy on large benchmark supervised labeled datasets compared to previous methods.
@@ 1 Introduction
Changepoint detection serves as a vital tool in numerous real-life applications by pinpointing significant transitions or sudden changes in data patterns, ranging from detecting market trends in finance (Lattanzi and Leonelli, 2021), monitoring disease outbreaks in healthcare (Muggeo and Adelfio, 2010), enhancing network security (Tartakovsky et al., 2013), monitoring environmental changes (Reeves et al., 2007) and more.
Several dynamic programming algorithms have been successfully applied to changepoint detection, including Optimal Partitioning (OPART) (Jackson et al., 2005), Functional Pruning Optimal Partitioning (FPOP) (Hocking et al., 2014), and Labeled Optimal Partitioning (LOPART) (Hocking and Srivastava, 2023). FPOP offers a faster implementation of OPART while producing the same results. LOPART extends OPART’s capabilities by leveraging predefined changepoint labels, indicating the expected number of changepoints within specific location ranges. When no such labels are defined, LOPART behaves identically to OPART. A critical aspect of these algorithms (be it OPART, FPOP, or LOPART) is the penalty parameter which is defined as $\lambda$ in the optimization problem below:
@@
Where $P$ is the possible changepoint locations set (locations in this set vary depending on whether the algorithm used is OPART/FPOP or LOPART). $l(m_{i},d_{i})$ is typically the negative log likelihood of the parameter $m_{i}$ given the value $d_{i}$; smaller values indicate a better fit. $I[m_{i}\neq m_{i+1}]$ is the indicator function, equaling 1 if there’s a changepoint (i.e., $m_{i}\neq m_{i+1}$ ) and 0 otherwise, $\lambda>0$ represents the penalty parameter.
@@
Previous methods, such as those employing Bayesian Information Criterion (BIC) (Schwarz, 1978), linear models (Hocking and Srivastava, 2023; Rigaill et al., 2013), ALPIN (Truong et al., 2017) and Maximum Margin Interval Trees (MMIT) (Drouin et al., 2017) have made good attempts to ascertain the optimal $\lambda$ value. However, these methods rely on using linear models or decision trees as learning models may constrain the ability to capture complex patterns.
@@
Figure 3: The loss value of two loss functions is determined by comparing the prediction to the target. In the left plot, the squared error loss function yields a value of 0 when the prediction $\hat{y}$ reaches the target point $y$. The right plot illustrates the squared hinge error (this loss function is utilized to predict values falling within a target interval rather than a single target point), where the loss value reaches 0 when the prediction $\hat{y}$ falls within the interval $[y_{l}+1,y_{u}-1]$.
@@ 2.1 Problem setting
Changepoint detection algorithms take into the sequence $\mathbf{d}$ along with the penalty parameter $\lambda$ as input and then generate the set of detected changepoints. From each sequence $\mathbf{d}$, $M$ distinct labels $\{(s_{i},e_{i},c_{i})\}_{i=1}^{M}$ emerge within the sequence, where $c_{i}$ denotes the expected number of changepoints between points at locations $s_{i}$ and $e_{i}$. Since ALPIN (Truong et al., 2017) requires a predefined expert sequence partition (or equivalently, each positive label having a size of 1, which is a specific special case), it is not suitable for this problem setting. There are two types of label errors: in each label, if the number of detected changepoints greater than $c_{i}$, it’s a false positive; if smaller, it’s a false negative. Finding the optimal set of detected changepoints require considering a range of values for the penalty parameter $\lambda$ (rather than just one specific value) that minimizes the total number of label errors. Consequently, the value of penalty parameter $\lambda$ to be predicted should also fall within an optimal interval, see Figure 2.
This type of problem, where the objective is to find values falling within an interval, is referred to as the interval regression problem. In the interval regression problem, there are four types of intervals: uncensored $(-\infty<y_{l}=y_{u}<\infty)$, censored $(-\infty<y_{l}<y_{u}<\infty)$, left-censored $(-\infty=y_{l}<y_{u}<\infty)$ and right-censored $(-\infty<y_{l}<y_{u}=\infty)$. According to the survey conducted by Wang et al. (2019), the majority of classic statistical parametric models (Tobin, 1958; Kalbfleisch and Prentice, 2002) are learned by assuming the data follow a particular theoretical distribution (commonly used distributions in parametric censored regression models are normal, exponential) and use maximum likelihood to estimate the parameters for these models. This model learning approach differs from the one that will be presented in this section, which involves learning the model by minimizing a specific loss function. While both model learning approaches yield very similar results, minimizing a specific loss function is an simpler approach and being easier to implement.
@@
As a result, squared error, which is commonly used as a loss function in point estimate regression, cannot be employed for this problem. This type of problem requires a different loss function. This loss function, known as squared hinge loss, is defined by Rigaill et al. (2013). The loss function resembles the squared error loss function, except it achieves a loss value of 0 within a target interval, as depicted in Figure 3, unlike the squared error loss function which only achieves 0 loss at a single target point.
@@
$l(\hat{y},y)=l(\hat{y},[y_{l},y_{u}])=\big{(}\text{ReLU}(y_{l}-\hat{y}+1)\big{)}^{2}+\big{(}\text{ReLU}(\hat{y}-y_{u}+1)\big{)}^{2}$
where $\hat{y}$ is the predicted value of penalty parameter $\lambda$ and $y=[y_{l},y_{u}]$, as $y_{l}<y_{u}$ is the optimal interval (lower, upper).
In summary, for each sequence, a vector of sequence features (e.g., length, variance, value range, $\dots$) and the optimal interval of penalty parameter $\lambda$ are obtained. During model training, the vector of sequence features is utilized as input to predict the value of penalty parameter $\lambda$, aiming for it to fall within the optimal interval by minimizing the squared hinge loss function.
Figure 4: Comparative Model Diagram: Previous Methods versus Our Proposal. From the raw sequence data, various sequence features are extracted such as length, mean, variance, standard deviation, minimum value, maximum value, quartiles, sum of absolute differences, $\dots$ and various transformation functions are applied such as log, absolute, square root, square, or combinations of these. This process yields a large vector of sequence features. However, only the features without missing data or non-infinity values are utilized for learning the penalty parameter value. While previous methods rely on BIC, linear models, or decision trees for penalty parameter prediction, our proposed approach adopts MLPs for enhanced performance.
@@ 2.2 Previous works
Following is a list of previous studies that have worked on predicting the value of $\lambda$, including the Bayes Information Criterion (BIC), linear models and decision trees. Define length as $N_{i}$ and variance as $\sigma_{i}$ for the $i^{th}$ sequence. To maintain consistency with the presentation of the $\lambda$ prediction problem in Rigaill et al. (2013) and Hocking and Srivastava (2023), the following sections will present the prediction problem as predicting $\log\lambda$ instead of $\lambda$.
@@
The Bayesian Information Criterion (BIC) proposed by Schwarz (1978) predicts $\log\lambda_{i}=\log\log N_{i}$ for each $i^{th}$ sequence. Notably, this model is unsupervised, it does not involve learning any parameters.
@@
This method constructs feature vectors $\mathbf{x}_{i}$ from $i^{th}$ sequence’s features. Prediction of $\log\lambda_{i}=\mathbf{x}_{i}^{T}w+b$ is performed, where parameters $w$ and $b$ are learned using convex optimization with a squared hinge loss. Hocking and Srivastava (2023) utilizes a single feature $\mathbf{x}_{i}=\log\log N_{i}$, while Rigaill et al. (2013) employs a feature vector $\mathbf{x}_{i}=[\log\log N_{i},\log\sigma_{i}]^{T}$. Moreover, Rigaill (2013) employs various statistical features and transformations (square, square root, absolute, log, loglog, $\dots$) to create a large feature vector $\mathbf{x}=[x_{1},x_{2},\dots]$, followed by the application of L1 regularization to address any redundant features.
MMIT
This method, introduced by Drouin et al. (2017), shares a similar concept with regression trees in Breiman et al. (1984). Instead of splitting based on minimizing squared error within regions like in Breiman et al. (1984), MMIT minimizes the hinge loss within each region. The optimal tree architecture, determined by hyperparameters such as maximum depth, minimum sample split and loss margin, is selected through cross-validation on the train set.
They are baseline algorithms will be used as a comparison for the new method. Their limitations and how they are addressed will be discussed in the following section.
Figure 5: Graphs illustrating the relationship between four features and target intervals. Variance and length are selected because previous studies have incorporated them into the model. Upon examination, we observe that the logarithms of range value, and sum of absolute differences exhibit a discernible increasing monotonic linear relationship with the target interval, indicating their potential as good features.
@@ 3 Novelty and contribution
Previous methods are constrained by a limitation: they rely on relatively simple models, such as linear models or decision trees. These models may not fully capture the complexity inherent in the underlying data, potentially leading to suboptimal outcomes.
Our method employs Multi-Layer Perceptrons (MLPs) to predict the value of $\lambda$. Leveraging MLPs for $\lambda$ prediction offers a distinct advantage, as they can extract pertinent hidden features from raw data, mitigating the need for manual feature engineering. Figure 4 provides a comparison summary between our proposed method and previous approaches. In MLPs, employing an excessive number of features is discouraged as the inclusion of irrelevant features can deteriorate the performance. To leverage MLPs effectively, we focus on selecting a subset of key features. In addition to the sequence length and variance utilized in previous methods, we incorporate two additional sequence features: the value range and the sum of absolute differences
@@
(length $N_{i}$ — variance $\sigma_{i}$ — value range $r_{i}$ — sum of absolute difference $s_{i}$) Model Features Model Regularization Citation BIC.1 $\log\log N_{i}$ unsupervised None Schwarz (1978) linear.1 $\log\log N_{i}$ linear None Hocking and Srivastava (2023) Rigaill et al. (2013) Lavielle (2005) linear.2 $\log\log N_{i},\log\sigma_{i}$ linear.4 $\log\log N_{i},\log\sigma_{i},\log r_{i},\log\log s_{i}$, linear.117 117 features linear L1 Rigaill et al. (2013) mmit.1 $N_{i}$ tree max depth min split sample loss margin Drouin et al. (2017) mmit.2 $N_{i},\sigma_{i}$ mmit.4 $N_{i},\sigma_{i},r_{i},s_{i}$, mmit.117 117 features mlp.1 $\log\log N_{i}$ MLP hidden layers number neurons per layer early stopping proposed mlp.2 $\log\log N_{i},\log\sigma_{i}$ mlp.4 $\log\log N_{i},\log\sigma_{i},\log r_{i},\log\log s_{i}$ mlp.117 117 features
@@ 4 Experiments
In this section, we will delve into the detailed implementation of our proposed approach aimed at enhancing reproducibility. We begin by processing raw sequence data, with given labels for each sequence, and extracting a set of features along with the target interval using the algorithm outlined by (Rigaill et al., 2013). Subsequently, we implement all baseline methods as well as our proposed method to generate predicted values for $\log\lambda$ (the summary of all employed models is provided in Table 1) Following this, we employ OPART with the predicted values of $\log\lambda$ to obtain the set of changepoints. Finally, leveraging both the set of changepoints and the set of labels, we compute the accuracy rate.
@@
This study employs three large datasets: two DNA copy number profiles sourced from neuroblastoma tumors (Rigaill et al., 2013) (available at https://github.com/tdhock/neuroblastoma-data), known for their detailed (3730 sequences) and systematic (3418 sequences) data collection. And the last one is a large epigenomic dataset comprising 17 sub-datasets (4913 sequences total) (Hocking et al., 2016) (accessible at https://archive.ics.uci.edu/ml/machine-learning-databases/00439/peak-detection-data.tar.xz).
@@
For each dataset, every sequence is assigned a unique identifier, referred to as sequenceID. The dataset is then divided into six folds based on sequenceID. During each iteration, one fold is designated as the test set, while the remaining five folds form the train set. This process is iterated six times, yielding six test accuracy rates. Figure 6 illustrates the median accuracy rate across all six test folds, along with the corresponding 25th and 75th percentile intervals.
@@
Below are presented the features, target interval, MLP configurations, as well as the loss function and optimizer
Features: Four sets of features are explored:
  • 1 feature: sequence length, same as Schwarz (1978) and Hocking and Srivastava (2023)
  • 2 features: sequence length and variance, same as Rigaill et al. (2013)
  • 117 features: same as Rigaill et al. (2013) and Drouin et al. (2017). There are 117 features selected from a large vector of 365 features, as only those without missing data (values too small for the computer to store) or infinity values are retained.
  • 4 features: sequence length, variance, value range, sum of absolute difference (a subset of 117 features to utilize MLPs)
@@
Figure 6: Median test accuracy and interquartile range (25% to 75%) across 6 folds for each method. To enhance clarity in the comparison visualization, method BIC and some small accuracies have been omitted (in the dataset detailed, methods mmit.1, mmit.2, and mmit.4 are excluded. In the dataset systematic, methods mmit.1 and mmit.2 are omitted. In the dataset epigenomic, method mmit.117 is not included). Across all three datasets, mlp.4 exhibits superior accuracy compared to all baseline methods, boasting higher median accuracy as well as improved accuracy interquartile.
@@
Target intervals: Every sequence within each dataset comes with its own set of labels, indicating the expected number of changepoints within specific regions. For each sequence, the target interval is the range of $\lambda$ value that minimizes the number of label errors (for each sequence with its set of detected changepoints, the minimum number of label errors is 0, and the maximum number of label errors is equal to the number of labels assigned to the sequence), which can be achieved by employing the algorithm described by Rigaill et al. (2013).
@@
Model Configurations Implementation: All baseline models (BIC, linear and decision tree) and MLPs with four different sets of features were employed (13 of them). To implement each MLP model, a two-fold cross-validation was employed on the train set to determine the optimal number of hidden layers and the number of neurons per hidden layer. This selection was based on achieving the highest accuracy rate on the corresponding validation set. Different MLP configurations were validated, including models with 1 to 4 hidden layers and all hidden layer of one model can have sizes of 2, 4, 8, 16, 32, 64, 128, 256 or 512 neurons.
@@
Number of iterations with Early Stopping: The early stopping technique, recommended by Prechelt (2012), controls the number of iterations. It uses the "patience" parameter to decide when to stop learning. During neural network training, setting a large maximum number of iterations and specifying the value for patience parameter allows for monitoring the training process. If the loss value from the train set does not decrease after the specified number of patience iterations, the training process is halted. This technique helps prevent overfitting. In our study, we set the maximum number of iterations to 12000 and patience parameter value to 20 iterations.
Random seed setting: Due to the use of random seeds, each time we initialize the parameters in an MLP model, we obtain a different set of initial values. Consequently, this leads to variability in the trained models. However, after testing with numerous random seeds, we observed that the outcomes closely resemble those depicted in Figure 6.
@@
Loss function and optimizer: The Adam optimizer is employed to minimize the squared hinge loss function (linear models and MLPs).

### v3

# Enhancing Changepoint Detection: Penalty Learning through Deep Learning Techniques

@@ Abstract
Changepoint detection, a technique for identifying significant shifts within data sequences, is crucial in various fields such as finance, genomics, medicine, etc. Dynamic programming changepoint detection algorithms are employed to identify the locations of changepoints within a sequence, which rely on a penalty parameter to regulate the number of changepoints. To estimate this penalty parameter, previous work uses simple models such as linear or tree-based models. This study introduces a novel deep learning method for predicting penalty parameters, leading to demonstrably improved changepoint detection accuracy on large benchmark supervised labeled datasets compared to previous methods.
keywords:
changepoint detection, supervised learning, deep learning, penalty parameter
@@ 1 Introduction
Changepoint detection serves as a vital tool in numerous real-life applications by pinpointing significant transitions or sudden changes in data patterns, ranging from detecting market trends in finance [1], monitoring disease outbreaks in healthcare [2], enhancing network security [3], monitoring environmental changes [4] and more.
To determine the optimal partitioning of a sequence given a fixed number of segments, several efficient algorithms can be employed to achieve precise results [5, 6, 7]. However, in many instances, the number of segments is not predetermined and must instead be inferred from the data. To address this challenge, several dynamic programming algorithms have been successfully applied to changepoint detection, including Optimal Partitioning (OPART) [8] with its variations such as Functional Pruning Optimal Partitioning (FPOP) [9] or Pruned Exact Linear Time (PELT) [7]. Labeled Optimal Partitioning (LOPART) [10] extends OPART’s capabilities by leveraging predefined changepoint labels, indicating the expected number of changepoints within specific location ranges. When no such labels are defined, LOPART behaves identically to OPART. A critical aspect of these algorithms (be it OPART or LOPART) is the penalty parameter which is defined as $\lambda$ in the optimization problem below:
@@
Where $P$ is the possible changepoint locations set (locations in this set vary depending on whether the algorithm used is OPART or LOPART). $l(m_{i},d_{i})$ is typically the negative log likelihood of the parameter $m_{i}$ given the value $d_{i}$; smaller values indicate a better fit. $I[m_{i}\neq m_{i+1}]$ is the indicator function, equaling 1 if there’s a changepoint (i.e., $m_{i}\neq m_{i+1}$ ) and 0 otherwise, $\lambda\geq 0$ represents the penalty parameter.
@@
Previous methods, such as those employing Bayesian Information Criterion (BIC) [11], linear models [10, 12], ALPIN [13], Maximum Margin Interval Trees (MMIT) [14] and Accelerated Failure Time (AFT) models in XGBoost [15] have made good attempts to ascertain the optimal $\lambda$ value. However, these methods rely on using linear or tree-based models as learning models may constrain the ability to capture complex patterns.
@@
Figure 3: The loss value of two loss functions is determined by comparing the prediction to the target. In the left plot, the squared error loss function yields a value of 0 when the prediction $\hat{y}$ reaches the target point $y$. The right plot illustrates the squared hinge error (this loss function is utilized to predict values falling within a target interval rather than a single target point), where the loss value reaches 0 when the prediction $\hat{y}$ falls within the interval $[y_{l}+\epsilon,y_{u}-\epsilon]$ where margin $\epsilon=1$.
@@ 2.1 Problem setting
Changepoint detection algorithms take into the sequence $\mathbf{d}$ along with the penalty parameter $\lambda$ as input and then generate the set of detected changepoints. From each sequence $\mathbf{d}$, $M$ distinct labels $\{(s_{i},e_{i},c_{i})\}_{i=1}^{M}$ emerge within the sequence, where $c_{i}$ denotes the expected number of changepoints between points at locations $s_{i}$ and $e_{i}$. Since ALPIN [13] requires a predefined expert sequence partition (alternatively, this corresponds to each label having a size of 1 having 1 changepoint, i.e., $s_{i}=e_{i}$ and $c_{i}=1$, which represents a specific special case), it is not suitable for this problem setting. There are two types of label errors: in each label, if the number of detected changepoints greater than $c_{i}$, it’s a false positive; if smaller, it’s a false negative.
Finding the optimal set of detected changepoints require considering a range of values for the penalty parameter $\lambda$ (rather than just one specific value) that minimizes the total number of label errors. Consequently, the value of penalty parameter $\lambda$ to be predicted should also fall within an optimal interval, see Figure 2. This type of problem, where the objective is to find values falling within an interval, is referred to as the interval regression. In the interval regression problem, there are four types of intervals: uncensored $(-\infty<y_{l}=y_{u}<\infty)$, interval-censored $(-\infty<y_{l}<y_{u}<\infty)$, left-censored $(-\infty=y_{l}<y_{u}<\infty)$ and right-censored $(-\infty<y_{l}<y_{u}=\infty)$.
Accelerated failure time (AFT) models, a class of linear models for censored outcomes, have been extensively studied [16]. In recent years, L1-regularized variants have been developed to handle high-dimensional datasets [17, 18]. Nonlinear methods for censored data, such as decision trees [19], Random Forests [20], and Support Vector Machines [21], have also been explored, though these techniques are limited to right-censored and uncensored data. In 2022, a variation of the AFT model, implemented in XGBoost, can handle various forms of censoring by applying an exponential transformation to ensure non-negative target intervals before fitting the model [15]. AFT models typically assume the data follow a specific distribution, such as normal, Weibull, exponential, etc, then use maximum likelihood estimation to learn the model parameters in parametric censored regression settings.
Another approach to the interval regression problem involves supervised machine learning, where the model is trained by minimizing a specific loss function. This includes methods such as the L1-regularized linear model [12] and the maximal margin interval tree (MMIT) [14].
While both model learning approaches (AFT models and supervised machine learning models) yield very similar results [15], minimizing a specific loss function offers a simpler, more intuitive approach. It does not require prior assumptions about the data distribution, which can be unclear to non-experts.
@@
As a result, squared error or absolute error, which is commonly used as a loss function in point estimate regression, cannot be employed for this problem. This type of problem requires a different loss function. This loss function, known as hinge loss, was first defined by Rigaill et al. [12], with its variants later introduced by Drouin et al. [14]. The loss function resembles the squared error (or absolute error) loss function, except it achieves a loss value of 0 within a target interval, as depicted in Figure 3, unlike the squared error (or absolute error) loss function which only achieves 0 loss at a single target point.
@@
$l(\hat{y},y)=l(\hat{y},[y_{l},y_{u}])=\big{(}\text{ReLU}(y_{l}-\hat{y}+\epsilon)\big{)}^{p}+\big{(}\text{ReLU}(\hat{y}-y_{u}+\epsilon)\big{)}^{p}$ (1)
In this context, $\epsilon\geq 0$ represents the margin length (Rigaill et al. [12] selected a fixed $\epsilon=1$, while Drouin et al. [14] treated $\epsilon$ as a hyperparameter, determined via cross-validation). The parameter $p$, which can be either 1 or 2, defines the type of loss function (Rigaill et al. [12] chose $p=2$, while Drouin et al. [14] considered both $p=1$ and $p=2$). The predicted value of the penalty parameter is denoted by $\hat{y}$, and $y=[y_{l},y_{u}]$ represents the optimal interval, where $y_{l}\leq y_{u}$ define the lower and upper bounds, respectively.
In summary, for each labeled training sequence, a vector of sequence features is derived (e.g., length, mean, minimum value, maximum value, variance, value range, …) and the optimal interval for the penalty parameter $\lambda$ is determined using Algorithm 1 in [12]. During model training, this feature vector serves as input to predict the value of $\lambda$, aiming for it to fall within the optimal interval by minimizing the squared hinge loss function. The predicted value of $\lambda$ is then used for partitioning new unlabeled sequences.
Figure 4: Comparative Model Diagram - Previous Methods versus Our Proposal. From the raw sequence data, various sequence features are extracted such as length, mean, variance, standard deviation, minimum value, maximum value, quartiles, sum of absolute differences, $\dots$ and various transformation functions are applied such as log, absolute, square root, square, or combinations of these. This process yields a large vector of sequence features. However, only the features without missing data or non-infinity values are utilized for learning the penalty parameter value. While previous methods rely on BIC, linear models, or decision trees for penalty parameter prediction, our proposed approach adopts MLPs for enhanced performance.
@@ 2.2 Previous works
Following is a list of previous studies that have worked on predicting the value of $\lambda$, including the Bayes Information Criterion (BIC), linear and tree-based models. Define length as $N_{i}$ and variance as $\sigma_{i}$ for the $i^{th}$ sequence. To maintain consistency with the presentation of the $\lambda$ prediction problem in [12] and [10], the following sections will present the prediction problem as predicting $\log\lambda$ instead of $\lambda$.
@@
The Bayesian Information Criterion (BIC) proposed by Schwarz [11] predicts $\log\lambda_{i}=\log\log N_{i}$ for each $i^{th}$ sequence. Notably, this model is unsupervised, it does not involve learning any parameters.
@@
This method constructs feature vectors $\mathbf{x}_{i}$ from $i^{th}$ sequence. The prediction model $\log\lambda_{i}=\mathbf{x}_{i}^{T}w+b$ is trained using one gradient descent technique named Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) algorithm by Chambolle and Dossal [22], which optimizes the parameters $w$ and $b$ through convex optimization with a squared hinge loss function. Hocking and Srivastava [10] utilizes a single feature $\mathbf{x}_{i}=\log\log N_{i}$, while Rigaill et al. [12] employs a feature vector $\mathbf{x}_{i}=[\log\log N_{i},\log\sigma_{i}]^{T}$. Moreover, Rigaill et al. [12] employs various statistical features and transformations (square, square root, absolute, log, loglog, $\dots$) in the vector of sequence features to create a large feature vector $\mathbf{x}=[x_{1},x_{2},\dots]$, followed by the application of L1 regularization to address any redundant features.
Maximum Margin Interval Trees (MMIT)
This method, introduced by Drouin et al. [14], shares a similar concept with regression trees by Breiman et al. [23]. Instead of splitting based on minimizing squared error within regions like in [23], MMIT minimizes the hinge loss within each region (can be the squared hinge loss presented in Figure 3 or absolute hinge loss). The optimal tree architecture, determined by hyperparameters such as maximum depth, minimum sample split and loss margin, is selected through cross-validation on the train set.
AFT Model in XGBoost
This model is a variation of the original AFT model [16], utilizing XGBoost to construct an ensemble of decision trees as the predictor. To adapt Accelerated Failure Time (AFT) models for gradient boosting frameworks like XGBoost, the model formulation is revised from the original AFT model as follows:
$\log(\lambda_{i})=\mathcal{T}(\mathbf{x}_{i})+\epsilon_{i}$
Here, $\mathcal{T}(\mathbf{x}_{i})$ represents the output from the decision tree ensemble based on input $\mathbf{x}_{i}$ (replacing the linear predictor $\mathbf{x}_{i}^{\top}\boldsymbol{\beta}$ in the original AFT model). The term $\epsilon_{i}$ is a random error following a specified distribution (e.g., normal, log-normal, Weibull).
To train this model, a distribution for the error term is selected, assuming that the model output is monotonically increasing with respect to the error. The likelihood is then constructed to accommodate various types of targets, and maximum likelihood estimation is used to train the model.
@@ 3 Novelty and contribution
Figure 5: Graphs illustrating the relationship between four features and target intervals. Variance and length are selected because previous studies have incorporated them into the model. Upon examination, we observe that the range value and sum of absolute differences exhibit a discernible increasing monotonic linear relationship with the target interval, indicating their potential as good features.
Previous methods are constrained by a limitation: they rely on relatively simple models, such as linear models or tree-based models. These models may not fully capture the complexity inherent in the underlying data, potentially leading to suboptimal outcomes.
Our method employs Multi-Layer Perceptrons (MLPs) to predict the value of $\lambda$. Leveraging MLPs for $\lambda$ prediction offers a distinct advantage, as they can extract pertinent hidden features from raw data, mitigating the need for manual feature engineering. Figure 4 provides a comparison summary between our proposed method and previous approaches. In MLPs, using an excessive number of features is not ideal, as the inclusion of noisy or irrelevant features can degrade performance To leverage MLPs effectively, we focus on selecting a subset of key features. In addition to the sequence length and variance utilized in previous methods, we incorporate two additional sequence features: the value range and the sum of absolute differences
@@
Model Features Model Type Regularization Citation
BIC.1 $\log\log N_{i}$ unsupervised None [11]
linear.1 $\log\log N_{i}$ linear None [10] [12] [24]
linear.2 $\log\log N_{i},\log\sigma_{i}$
linear.4 $\log\log N_{i},\log\sigma_{i},\log r_{i},\log\log s_{i}$
linear.all all features L1 [12]
mmit.1 $N_{i}$ tree max depth [14]
mmit.2 $N_{i},\sigma_{i}$ min split sample
mmit.4 $N_{i},\sigma_{i},r_{i},s_{i}$ loss margin
mmit.all all features
aft_xgboost.1 $N_{i}$ ensemble tree [15]
aft_xgboost.2 $N_{i},\sigma_{i}$ learning rate
aft_xgboost.4 $N_{i},\sigma_{i},r_{i},s_{i}$ max depth
aft_xgboost.all all features
mlp.1 $\log\log N_{i}$ MLP hidden layers number proposed
mlp.2 $\log\log N_{i},\log\sigma_{i}$ neurons per layer
mlp.4 $\log\log N_{i},\log\sigma_{i},\log r_{i},\log\log s_{i}$ early stopping
mlp.all all features
[…]
00footnotetext: Notations: length $N_{i}$ — variance $\sigma_{i}$ — value range $r_{i}$ — sum of absolute difference $s_{i}$
@@ 4 Experiments
In this section, we will delve into the detailed implementation of our proposed approach aimed at enhancing reproducibility. We begin by processing raw sequence data, with given labels for each sequence, and extracting a set of features along with the target interval using the algorithm outlined by Rigaill et al. [12]. Subsequently, we implement all baseline methods as well as our proposed method to generate predicted values for $\log\lambda$ (the summary of all employed models is provided in Table 1). Following this, we employ OPART with the predicted values of $\log\lambda$ to obtain the set of changepoints. Finally, leveraging both the set of changepoints and the set of labels, we compute the accuracy rate.
@@
This study employs three large datasets: two DNA copy number profiles sourced from neuroblastoma tumors [12] (available at https://github.com/tdhock/neuroblastoma-data), known for detailed (3730 sequences) and systematic (3418 sequences) data collection. And the last one is a large epigenomic dataset comprising 17 sub-datasets (4913 sequences total) [25] (accessible at https://archive.ics.uci.edu/ml/machine-learning-databases/00439/peak-detection-data.tar.xz).
@@
For each dataset, every sequence is assigned a unique identifier, referred to as sequenceID. The dataset is then divided into six folds based on sequenceID. During each iteration, one fold is designated as the test set, while the remaining five folds form the train set. Therefore this process is iterated six times, yielding six test accuracy rates. Figure 6 illustrates the mean accuracy rate across all six test folds, along with the corresponding plus-minus one standard deviation intervals.
Figure 6: Mean test accuracy and ±1 standard deviation across 6 folds for each method. To enhance clarity in the comparison visualization, method BIC and some small accuracies have been omitted (in the dataset detailed, methods mmit.1, mmit.2, and mmit.4 are excluded; in the dataset systematic, methods mmit.1 and mmit.2 are omitted; in the dataset epigenomic, method mmit.all is not included). Across all three datasets, mlp.4 exhibits superior accuracy compared to all baseline methods, boasting higher mean accuracy as well as improved accuracy variability.
@@
Below are presented the features, target interval, baseline models implementation, MLP configurations, as well as the loss function and optimizer.
Features
Four sets of features are explored:
  • 1 feature: sequence length, same as [11] and [10]
  • 2 features: sequence length and variance, same as [12]
  • 4 features: sequence length, variance, value range, sum of absolute difference (a subset of all features to utilize MLPs)
  • all features: same as [12] and [14]. Features are selected from a large vector, as only those without missing data (values too small for the computer to store), infinity values are retained, or zero variance.
@@
Target intervals
@@
Every sequence within each dataset comes with its own set of labels, indicating the expected number of changepoints within specific regions. For each sequence, the target interval is the range of $\lambda$ value that minimizes the number of label errors (for each sequence with its set of detected changepoints, the minimum number of label errors is 0, and the maximum number of label errors is equal to the number of labels assigned to the sequence).
Baseline Models Implementation
All baseline models (BIC, linear, MMIT and AFT model in XGBoost) with four different sets of features were employed.
For the linear models, we utilize the R package penaltyLearning [26]. In the case without regularization, we apply the Max Margin Interval Regression model using corresponding features. For L1 regularization, we use the same Max Margin Interval Regression model with L1 regularization. The L1 regularization parameter begins at 0.001 and is increased by a factor of 1.2 until no features remain. Cross-validation is employed to determine the optimal L1 regularization value. In addition to using the R package, we also implemented this model in Python. Instead of using FISTA, we employed the widely-used Adam optimizer for gradient descent and achieved same results to those obtained with the author’s original code.
The MMIT method is implemented using the R package mmit [27]. We use cross-validation with cv = 2 to select the best hyperparameters. The hyperparameters considered include a list for max_depth values (0, 1, 5, 10, 20, Inf), a list for margin (or $\epsilon$ from equation 1) values (0, 1, 2), a list for loss_type values (hinge, square – equivalent to $p=1$ or $p=2$ in Equation 1) and a list for min_sample values (0, 1, 2, 4, 8, 16, 20).
The AFT model is implemented using the Python package xgboost. We use cross-validation with cv=2 to select the best hyperparameters. The hyperparameters considered include a list of learning_rate values (0.01, 0.1, 0.2) and a list of max_depth values (6, 8, 10, 20). Training data is prepared by converting feature and target data, with the target values transformed using the exponential function to ensure non-negativity. The selected hyperparameters from cross-validation step are then used to train the final model on the entire training dataset, with a maximum of 10,000 boosting rounds and early stopping after 100 rounds without improvement. Predictions are made on the test dataset and transformed back to the original scale using a logarithm.
MLP Configurations
@@
To implement each MLP model, a two-fold cross-validation was employed on the train set to determine the optimal number of hidden layers and the number of neurons per hidden layer. This selection was based on achieving the highest accuracy rate on the corresponding validation set. Different MLP configurations were validated, including models with 1 to 4 hidden layers and all hidden layer of one model can have sizes of 2, 4, 8, 16, 32, 64, 128, 256 or 512 neurons.
MLP Implementation
All MLP models were trained using PyTorch, with ReLU activation functions applied to all hidden layers to provide greater generalization compared to the linear model.
Number of iterations with Early Stopping
@@
There is one early stopping technique, presented by Prechelt [28], controls the number of iterations. It uses the “patience” parameter to decide when to stop learning. During neural network training, setting a large maximum number of iterations and specifying the value for patience parameter allows for monitoring the training process. If the loss value from the train set does not decrease after the specified number of patience iterations, the training process is halted. This technique helps prevent overfitting. In our study, we set the maximum number of iterations to 12000 and patience parameter value to 20 iterations.
Loss function and optimizer
@@
The Adam optimizer is employed to minimize the squared hinge loss function with margin $\epsilon=1$ (linear models and MLPs).